the cost of a renting car in 2 500 pesos day plus 20 pesos per kilometer travelled. find the cost of the renting car that will be used for a 2 day, 150 kilometer trip?

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

it's 2500 per day plus 20 per kilometer.
the car was rented for 2 days and went 150 kilometers.
the cost would be 5000 + 3000 = 8000 pesos.
the equation is c = 2500 * d + 20 * k
c is the total cost
d = the number of days
k = the number of kilometers.
when d = 2 and k = 150, the equation becomes:
c = 2500 * 2 + 20 * 150.
solve for c to get:
c = 5000 + 3000 = 8000 pesos



Answer by MathTherapy(10552)